Solve dy/dt + ycotx = e^cosx for x= pi by two and y= -2
1
Expert's answer
2020-03-25T10:59:04-0400

Mechanics | Relativity

We need to solve the given Differential equation

Solution:

Given equation is the Linear differential in the form "\frac {dy}{dx} + P(x) y = Q(x)"


Now compare the given equation "\frac {dy}{dx} + y \space cot x = e^ {cos \space x}" with the above form



"P(x) = cot \space x"

"Q(x) = e^{cos \space x}"

Integrating factor = I.F = "e^{\int P(x) dx} = e^ {\int cot x \space dx} = e^ {ln \space sin x} = sin \space x"


The solution of the Linear equation is,



"y \times (I.F) = \int (I.F) \times Q(x) dx + c"

"y \times sin x = \int sin x \space e^{cos x} dx + c"

"y \space sin x = - \int d(cos x) \space e^{cos x} dx + c"

"y \space sin x = - e^{cos x} + c"

Now we need to find the c value by plugging "x = \frac {\pi}{2} \space and \space y =-2"



"-2 (sin \frac{\pi}{2} ) = - e^{cos {\frac {\pi}{2 }}} + c"

"-2 =- e^0 + c \\
-2 = -1 +c \\
c = -1"

Finally the solution is ,

"y \space sin x = - e^{cos x} -1"